# Designed self-assembly of programmable colloidal atom-electron equivalents

### v3

@@ 2.2 EE-mediated interactions
In typical experimental PAE-EE systems, van der Waals interactions and electrostatic interactions can be neglected [30, 31]. Besides, we also neglect the depletion interaction originating from the EEs acting as depletants, particularly when EEs are sufficiently small. Based on the fact that PAEs are much larger than EEs [18], we assume that the motion of PAEs and EEs can be decoupled, and EEs move within the static external potential generated by PAEs and provide effective interactions among PAEs. Consequently, the effective interaction among PAEs can be expressed by
@@
where $M_{i}$ and $Q_{ij}$ are the total numbers of EEs in the adsorbed state $\mathcal{A}_{i}$ and bridging state $\mathcal{B}_{ij}$, respectively, and the constant indicates the reference to different configurations of PAEs. Similarly, the free energy of EEs for isolated PAE configurations $\beta F_{\rm att}^{\prime}$ is derived in Supplementary Material Sec. II.C. We notice that Eq. 6 can be numerically solved, and $M_{i}$ and $Q_{ij}$ in Eq. 9 are in closed forms for arbitrary $\lambda$ (see Supplementary Material Sec. II.B [32]). With the system consisting of $N$ particles, we obtain a set of $N$-element self-consistent equation with variables $(\bar{n}_{1},\bar{n}_{2},\cdots,\bar{n}_{N})$ from Eq. 6, which can be numerically solved by fixed-point iteration. In Supplementary Material Sec. II.F, we prove that there exists a unique solution $\{\bar{n}_{i}\}$ for $0\leq\bar{n}_{i}\leq n_{i}$ [32]).
The theory is based on the mean field approximation, assuming negligible fluctuations in EE occupancy and rapid DNA binding/unbinding kinetics compared to colloidal dynamics. This approximation holds well when many EEs participate, do not strongly interfere with each other, and binding equilibria can quickly adjust to changes in colloidal configurations. We note that when the EE is heterogeneously bivalent, i.e., $\lambda=1$, the free energy can be simplified as
$\beta F_{\rm att}=\sum_{i}\left(n_{i}\log\frac{\bar{n}_{i}}{n_{i}}+\sum_{j<i}Q_{ij}\right),$ (10)

### v4

@@ 2.2 EE-mediated interactions
In typical experimental PAE-EE systems, van der Waals interactions and electrostatic interactions can be neglected [30, 31]. Besides, we also neglect the depletion interaction originating from the EEs acting as depletants, particularly when EEs are sufficiently small. Detailed justifications for these approximations are provided in the Supplementary Material Sec. V.B [32]. Based on the fact that PAEs are much larger than EEs [18], we assume that the motion of PAEs and EEs can be decoupled, and EEs move within the static external potential generated by PAEs and provide effective interactions among PAEs. Consequently, the effective interaction among PAEs can be expressed by
@@
where $M_{i}$ and $Q_{ij}$ are the total numbers of EEs in the adsorbed state $\mathcal{A}_{i}$ and bridging state $\mathcal{B}_{ij}$, respectively, and the constant indicates the reference to different configurations of PAEs. Similarly, the free energy of EEs for isolated PAE configurations $\beta F_{\rm att}^{\prime}$ is derived in Supplementary Material Sec. II.C. We notice that Eq. 6 can be numerically solved, and $M_{i}$ and $Q_{ij}$ in Eq. 9 are in closed forms for arbitrary $\lambda$ (see Supplementary Material Sec. II.B [32]). With the system consisting of $N$ particles, we obtain a set of $N$-element self-consistent equation with variables $(\bar{n}_{1},\bar{n}_{2},\cdots,\bar{n}_{N})$ from Eq. 6, which can be numerically solved by fixed-point iteration. In Supplementary Material Sec. II.F, we prove that there exists a unique solution $\{\bar{n}_{i}\}$ for $0\leq\bar{n}_{i}\leq n_{i}$ [32]). We note that when the EE is heterogeneously bivalent, i.e., $\lambda=1$, the free energy can be simplified as
$\beta F_{\rm att}=\sum_{i}\left[n_{i}\log\frac{\bar{n}_{i}}{n_{i}}+\sum_{j<i}Q_{ij}\right],$ (10)
@@
The theory is based on the mean-field approximation, assuming negligible fluctuations in EE occupancy and rapid DNA binding/unbinding kinetics compared to colloidal dynamics. This approximation holds well when many EEs participate, do not strongly interfere with each other, and binding equilibria can quickly adjust to changes in colloidal configurations. Experimentally, EEs in assemblies are highly mobile, often exhibiting “metallic” behavior analogous to electrons in metals, where EEs roam freely throughout the colloidal crystal and reach equilibrium quickly [18, 19]. This high mobility underlies the validity of the mean-field approach in our parameter regimes, as fluctuations are efficiently averaged out and correlations remain weak. Even in the transition to more ionic-like structures upon crystallization, the delocalization-to-localization of EEs occurs continuously rather than abruptly [21], suggesting that the mean-field description remains appropriate until EEs become strongly localized or comparable in size to the PAE unit cell. Moreover, we do not explicitly include crowding among EE strands, as their concentrations are low and each EE typically carries only a few DNA arms under experimental conditions, making the steric penalty much smaller than for densely grafted PAE receptors. If necessary, excluded volume effects between EEs could be incorporated via an additional chemical-potential correction to the free energy, and this would introduce extra parameters without affecting the essential physics of multivalency  [33].